Solve.
1) y squared + 3 = 21

2) (x+5)squared = 16

Answer by gems(12)

1. y sqrd + 3 =21
y sqrd = 21-3
y sqrd = 18
y=sqrt 18
y= sqrt (9*2)
y= 3 sqrt 2
2. (x+5) sqrd = 16
x+5= sqrt 16
x+5 = +/- 4
x1=4-5
x1=-1
x2=-4-5
x2=-9